# Improving the Validity of Automatically Generated Feedback via Reinforcement Learning

### v1

@@ 2.1 Feedback Generation
There are many existing approaches for automatic feedback generation. One common method is to use engineered features to detect errors in student responses, and then use a rule-based system to provide relevant feedback or hints [3, 12, 15, 23, 27, 28]. This method is popular since it is interpretable and reliable but requires significant human effort to adapt to new question types. A recent and more general approach to feedback generation is using large language models (LLMs), either through prompting [1, 19, 21, 30, 33] or fine-tuning [10]. However, prompting pre-trained LLMs requires the existing models to have good behavior and understanding of educational goals, but fine-tuning can yield poor results without significant amounts of aligned training data. We address these concerns in our work by fine-tuning on an augmented dataset annotated with alignment labels.
@@ 2.2 Feedback Evaluation
Several recent works have used rubrics to evaluate feedback [10, 30, 33], and works in other domains have found success in using LLMs to evaluate open-ended text where their judgements correlate with human judgements [6, 13, 20]. However, most prior works on feedback generation tend to rely on human annotators for reliable evaluation [1, 9, 10, 30, 33]. One recent work [11] uses GPT-4 to evaluate math feedback with a rubric and finds high agreement with human annotations. However, they only use GPT-4 to evaluate human-written feedback, while we evaluate feedback written by both humans and LLMs. Including this LLM feedback helps us uncover GPT-4’s shortcomings in feedback evaluation, particularly that it can struggle to identify when feedback inaccurately addresses student errors or provides invalid suggestions.
@@ 3.1 Feedback Evaluation
  • Diagnositic (DIA.) The feedback correctly points out the error the student made or the misconception underlying their answer.
@@
We now define the rubric function, which assigns labels to any feedback, given a corresponding question and incorrect answer:
@@
We also define a final scalar-valued rubric score for a feedback message by aggregating the labels in $\mathbf{y}$, indicating the feedback’s overall quality:
@@
System
You are a math education expert.
Prompt
Your job is to evaluate feedback given to students on math problems.
@@ 5. Does the feedback have a positive or encouraging tone?
Output
1. Explanation: The feedback is correct in stating that the student has simplified, but not fully. The student stopped at $\frac{18}{84}$, which can be further simplified to $\frac{3}{14}$. So, the feedback does not make any incorrect statements. Answer: No.
2. Explanation: The feedback does not provide the correct answer. It only indicates that the student’s answer is not fully simplified. Answer: No.
3. Explanation: The feedback does not provide specific suggestions on how to improve the answer. It only states that the student has not fully simplified the fraction. Answer: No.
4. Explanation: The feedback correctly identifies that the student’s misconception is not fully simplifying the fraction. Answer: Yes.
5. Explanation: The feedback is neutral in tone. It points out the error but does not use any positive or encouraging language. Answer: No.
Result
$y^{\prime}_{\text{C}}=1$, $y^{\prime}_{\text{R}}=1$, $y^{\prime}_{\text{S}}=0$, $y^{\prime}_{\text{D}}=1$, $y^{\prime}_{\text{P}}=0$, $s^{\prime}=0.6$
@@ 3.2 Data Augmentation
We now detail our method for constructing an augmented dataset, which will be used for RL training as well as calculating agreement between GPT-4 and human annotations. For both of these tasks, we require both positive examples, i.e., feedback messages that score highly on the rubric, and negative examples, i.e., feedback that score poorly on the rubric. We observe that while most of the gold feedback messages in $\mathcal{D}$ score highly, feedback generated by prompting LLMs can be flawed or unaligned with the evaluation metric, thus scoring poorly. Therefore, we collect 3 LLM-augmented versions of $\mathcal{D}$, where each feedback $f^{(i)}_{j}$ is replaced with a generated version: $\mathcal{D}_{R}$, where feedback is generated using few-shot prompting with random in-context examples, $\mathcal{D}_{S}$, where feedback is generated using few-shot prompting with the most similar examples, and $\mathcal{D}_{Z}$, where feedback is generated using zero-shot prompting. We refer to the union of the original dataset and LLM-augmented datasets $\mathcal{D}^{\prime}=\bigcup\{\mathcal{D},\mathcal{D}_{R},\mathcal{D}_{S},\mathcal{D}_{Z}\}$ as the augmented dataset. We use GPT-4 to annotate the feedback messages in the augmented dataset, and describe how we use these annotations for RL training in the next section.
@@ 4.1 Dataset
We validate our framework using a dataset of middle school-level math multiple choice questions from a math learning platform. The questions cover a variety of number sense concepts including fractions, exponents, rounding, and many others. All questions and feedback messages are written by real math teachers, deployed to real students, and are generally high quality. There are a total of 1,956 questions in the dataset and each question has a total of 3 incorrect options and a ground truth human-written feedback for each. We remove questions that require images and ones with processing errors, resulting in 1,418 questions. We divide these into a train/validation/test split of 850/284/284 questions and correspondingly 2,550/852/852 incorrect answers and corresponding feedback.
@@ 4.2 Experimental Setting
We primarily use the instruction-tuned Llama-2 7B Chat model [32] from HuggingFace [34] for feedback generation, loaded with 8-bit quantization [7]. For both supervised fine-tuning (SFT) and DPO, we train LoRA adapters [8] on all weight matrices, setting $r=32$, $\alpha=16$, and dropout$=0.05$. We train using the AdamW optimizer with a learning rate of 3e-4 with warmup for 10% of steps and an effective batch size of 64 using gradient accumulation. We train for 3 epochs, which we find minimizes the loss on the validation set. For DPO, we initialize $\theta$ to the SFT weights, use the SFT model as the reference model, and set $\beta=0.5$. At inference time, we use greedy decoding and set the maximum new tokens to 200.
@@
When evaluating feedback, we report the average of each rubric label in $\mathbf{y}^{\prime}$ and the corresponding scores $s^{\prime}$ assigned by GPT-4. We note that GPT-4 will very rarely fail to assign labels when feedback is unrelated to the current question, in which case we automatically assign label values of 0. We use a temperature of 0 and 300 maximum new tokens for GPT-4 decoding. We additionally use two popular reference-based metrics where we use the human-written feedback as a gold reference: ROUGE-L (ROU.) [17] which is based on textual overlap, and the F1 of the BERTScore (BER.) [35] using the recommended microsoft/deberta-xlarge-mnli model, which is based on token-level semantic similarity.
@@ 4.3 Feedback Generation
We now show that we can improve both the correctness and alignment of generated feedback using our framework. We primarily focus on using Llama 2 Chat, an open-source LLM with 7B parameters, where we compare several versions of the model: Zero-Shot, i.e., simply prompting the base LLM, SFT, i.e., fine-tuning the base LLM on the gold feedback set, DPO (Score), i.e., training the base LLM with DPO only on the augmented dataset, and DPO (Score + Mismatch), i.e., training the base LLM with DPO on the augmented dataset and mismatched feedback. We additionally compare with the gold, human-written feedback in the dataset, as well as feedback generated by GPT-4. We use the same prompt for all methods, where we instruct the model to generate short and helpful feedback and to follow a version of the evaluation rubric.
@@ 4.3.1 Quantitative Analysis
Table 2 shows the average rubric labels and scores assigned by GPT-4 on all feedback in the test set, as well as the ROUGE-L and BERTScore values for reference. We see that DPO (Score + Mismatch) significantly improves the feedback scores compared to baselines (a 45% increase compared to Zero-Shot and SFT), showing that our data augmentation and training setup is highly effective at improving the quality of feedback generation with Llama 2. We additionally observe that including the mismatched feedback messages substantially increases the correctness of generated feedback, confirming their effectiveness as hard negative examples. Surprisingly, SFT does not outperform Zero-Shot on score, which shows that the standard fine-tuning setup is not effective for the specific task of feedback generation. We can also see that ROUGE-L and BERTScore are unreliable estimates of feedback quality since they are highest on SFT, primarily because it copies the style of the gold feedback the closest.
@@
We also performed qualitative studies to compare the outputs of the different methods and find cases where they succeed or fail; the main findings are:
  • Zero-Shot produces feedback with the right style but struggles to follow instructions, particularly by not identifying the error or revealing the correct answer, and is prone to hallucinations and numerical errors.
  • SFT produces feedback that is generally short and blunt, and usually attempts to identify the error although is often incorrect.
  • DPO (Score) produces feedback that attempts to identify the error, adding details and questions to provide implicit suggestions and increase positivity. It also produces incorrect outputs although less so than SFT.
  • DPO (Score + Mismatch) is more accurate than DPO (Score) in identifying the error.
  • GPT-4 produces feedback with smooth and coherent language but tends to avoid mistakes by not clearly pointing out the error. When it does detail the error, it is often incorrect, although less so than the other methods.
  • For all methods, most failures occur because the LLM cannot correctly identify the error in the incorrect answer. This failure happens more often when multiple reasoning steps are required to arrive at the student’s answer. This observation suggests that although state-of-the-art LLMs claim to excel at mathematical reasoning, they cannot conceive how students may make errors [11, 18, 19]. It may be necessary to prompt the LLM to perform necessary calculations before writing the feedback [29] and use calculators or code execution [5] when doing so.
To provide more specific examples of several of our findings, we show a representative question and incorrect answer along with corresponding feedback generated by each method in Table 3. In this example, we see that the human-written feedback directly points out the error in the incorrect answer that the student only considered the denominators, and also suggests simplifying the fraction on the right. GPT-4 prefers to be vague to avoid mistakes by pointing out that the ratio should be simplified and gives a suggestion based on this, but does not explicitly identify why the student’s answer was wrong. Zero-Shot does not point out the student’s error and proceeds to answer the question, even though it has been instructed to not reveal the answer; it then hallucinates an incorrect and irrelevant multiplication before finishing with a positive message. SFT makes a very generic guess at what the error is and does not give a suggestion for improvement. DPO (Score) guesses a more specific but incorrect error, gives a somewhat misleading suggestion, and uses questions to encourage the student. Finally, DPO (Score + Mismatch) correctly identifies the error and gives a helpful suggestion.
@@
Human
Not quite, the denominator has doubled but the fractions don’t have that relationship. Try simplifying 6/8 and then compare it to 1/4
GPT-4
Good effort, but remember to simplify the ratios first. Erica’s share is not twice as much, but a fraction of Rebecca’s. Try to compare 1/4 to 3/4 again. Keep going! ✓
Zero-shot
Great effort! You are on the right track. Since the ratio is 1/4: 3/4, Erica has 1/4, which is a third of 3/4. To find the number of marbles Erica has, you can multiply 3/4 by 3, which gives you 9. So, Erica has 9 marbles. Keep practicing, and you’ll get it right next time!
SFT
I think you have simplified the second fraction incorrectly.
DPO (Score)
I think you have simplified 6/8 to 1/2. If we simplify 6/8 to 3/4, how many marbles does Erica have? How many does Rebecca have? What is the difference between them?
DPO (Score + Mismatch)
I think you have found 8/4 = 2. But we need to simplify 6/8 to get the same units as 1/4. Then we can see how many more or less Erica has.
@@ 4.4 Feedback Evaluation
COR. 0.78 0.78 0.85 0.81
REV. 0.96 1.00 0.96 0.98
SUG. 0.75 0.67 0.79 0.72
DIA. 0.69 0.58 0.85 0.69
POS. 0.82 0.78 0.37 0.50
Avg. 0.80 0.76 0.76 0.74
Since we use GPT-4 to quantitatively evaluate feedback, we need to verify that GPT-4 can indeed label feedback accurately using our rubric. To do so, we randomly sample 80 feedback messages from the augmented test set and manually evaluate them on the rubric. To reduce bias, we do not show the human annotator GPT-4’s annotations or tell them whether a feedback message is human-written or LLM-generated. We measure the accuracy (Acc.), precision (Prec.), recall (Rec.), and F1 of GPT-4’s annotations with respect to human annotations on this set, and report the results in Table 4.
We observe that GPT-4 generally agrees with our annotations, with an average accuracy of 80% across labels. We also compute the Pearson correlation coefficient of the final rubric scores between GPT-4 and our annotations, resulting in 0.58, indicating moderate overall correlation. We note that ROUGE-L and BERTScore have correlations of 0.49 and 0.50 with our rubric scores, respectively. Not only are these numbers lower, but we note that they are biased upward since the human-written feedback messages, which generally have high rubric scores, automatically get ROUGE-L and BERTScore values of 1.
@@
COR. (FP)
Question: a jacket is reduced in a sale by $20\%$. a customer has a loyalty card for a further $10\%$ off the price. what is the total percentage reduction in the price of the jacket?
@@
COR. (FN)
Question: i am thinking of a number. i divide it by 0.5 and my answer is 10. what number am i thinking of?
@@
DIA. (FP)
Question: which fraction is equivalent $\frac{3}{7}$

### v2

@@ Abstract
Published in AIED 2024: The 25th International Conference on Artificial Intelligence in Education.
@@ 2.1 Feedback Generation
There are many existing approaches for automatic feedback generation. One common method is to use engineered features to detect errors in student responses, and then use a rule-based system to provide relevant feedback or hints [3, 12, 15, 25, 29, 30]. This method is popular since it is interpretable and reliable but requires significant human effort to adapt to new question types. A recent and more general approach to feedback generation is using large language models (LLMs), either through prompting [1, 20, 23, 32] or fine-tuning [10]. However, prompting pre-trained LLMs requires them to be capable of understanding educational goals, but fine-tuning can yield poor results without significant amounts of aligned training data. We address these concerns in our work by fine-tuning on an augmented dataset annotated with alignment labels.
@@ 2.2 Feedback Evaluation
Several recent works have used rubrics to evaluate feedback [10, 32], and works in other domains have found success in using LLMs to evaluate open-ended text where their judgements correlate with human judgements [6, 13, 22]. However, most prior works on feedback generation tend to rely on human annotators for reliable evaluation [1, 9, 10, 32]. One recent work [11] uses GPT-4 to evaluate math feedback with a rubric and finds high agreement with human annotations. However, they only use GPT-4 to evaluate human-written feedback, while we evaluate feedback written by both humans and LLMs. Including this LLM feedback helps us uncover GPT-4’s shortcomings in feedback evaluation, particularly that it can struggle to identify when feedback inaccurately addresses student errors or provides invalid suggestions.
@@ 3.1 Feedback Evaluation
  • Diagnostic (DIA.) The feedback correctly points out the error the student made or the misconception underlying their answer.
@@
We now define the rubric function, $\operatorname{r}$, which assigns labels to any feedback given a corresponding question and incorrect answer, and a final scalar-valued rubric score, $s$, which indicates the feedback’s overall quality:
@@
System You are a math education expert.
Prompt Your job is to evaluate feedback given to students on math problems.
@@ 5. Does the feedback have a positive or encouraging tone?
Output 1. Explanation: The feedback is correct in stating that the student has simplified, but not fully. The student stopped at $\frac{18}{84}$, which can be further simplified to $\frac{3}{14}$. So, the feedback does not make any incorrect statements. Answer: No. 2. Explanation: The feedback does not provide the correct answer. It only indicates that the student’s answer is not fully simplified. Answer: No. 3. Explanation: The feedback does not provide specific suggestions on how to improve the answer. It only states that the student has not fully simplified the fraction. Answer: No. 4. Explanation: The feedback correctly identifies that the student’s misconception is not fully simplifying the fraction. Answer: Yes. 5. Explanation: The feedback is neutral in tone. It points out the error but does not use any positive or encouraging language. Answer: No.
Result ($y^{\prime}_{\text{C}}=1$, $y^{\prime}_{\text{R}}=1$, $y^{\prime}_{\text{S}}=0$, $y^{\prime}_{\text{D}}=1$, $y^{\prime}_{\text{P}}=0$) $\rightarrow$ $s^{\prime}=0.6$
@@ 3.2 Data Augmentation
We now detail our method for constructing an augmented dataset, which will be used for RL training as well as calculating agreement between GPT-4 and human annotations. For both of these tasks, we require both positive examples, i.e., feedback messages that score highly on the rubric, and negative examples, i.e., feedback that score poorly on the rubric. We observe that while most of the gold feedback messages in $\mathcal{D}$ score highly, feedback generated by prompting LLMs can be flawed or unaligned with the evaluation metric, thus scoring poorly. Therefore, we collect 3 LLM-augmented versions of $\mathcal{D}$, where each feedback $f^{(i)}_{j}$ is replaced with a generated version: $\mathcal{D}_{R}$, where feedback is generated using few-shot prompting with random in-context examples, $\mathcal{D}_{S}$, where feedback is generated using few-shot prompting with the most similar examples, and $\mathcal{D}_{Z}$, where feedback is generated using zero-shot prompting. We refer to the union of the original dataset and LLM-augmented data $\mathcal{D}^{\prime}=\bigcup\{\mathcal{D},\mathcal{D}_{R},\mathcal{D}_{S},\mathcal{D}_{Z}\}$ as the augmented dataset. We use GPT-4 to annotate the feedback messages in this set, and we detail how we use these annotations for training in the next section.
@@ 4.1 Dataset
We validate our framework using a dataset of middle school-level math multiple choice questions from Eedi, a math learning platform. The questions cover a variety of number sense concepts including fractions, exponents, rounding, and many others. All questions and feedback messages are written by real math teachers, deployed to real students, and are generally high quality. There are a total of 1,956 questions in the dataset and each question has a total of 3 incorrect options and a ground truth human-written feedback for each. We remove questions that require images and ones with processing errors, resulting in 1,418 questions. We divide these into a train/validation/test split of 850/284/284 questions and correspondingly 2,550/852/852 incorrect answers and corresponding feedback.
@@ 4.2 Experimental Setting
We primarily use the instruction-tuned Llama-2 7B Chat model [34] from HuggingFace [35] for feedback generation, loaded with 8-bit quantization [7]. For both supervised fine-tuning (SFT) and DPO, we train LoRA adapters [8] on all weight matrices, setting $r=32$, $\alpha=16$, and dropout$=0.05$. We train using the AdamW optimizer with a learning rate of 3e-4 with warmup for 10% of steps and an effective batch size of 64 using gradient accumulation. We train for 3 epochs, which we find minimizes the loss on the validation set. For DPO, we set $\beta=0.5$ and use the SFT model for initialization and as the reference model, which empirically outperformed using the base model. At inference time, we use greedy decoding and set the maximum new tokens to 200.
@@
When evaluating feedback, we report the average of each rubric label in $\mathbf{y}^{\prime}$ and the corresponding scores $s^{\prime}$ assigned by GPT-4. We note that GPT-4 will very rarely fail to assign labels when feedback is unrelated to the current question, in which case we automatically assign label values of 0. We use a temperature of 0 and 300 maximum tokens for GPT-4 decoding. We also use two popular reference-based metrics with the human-written feedback as reference: ROUGE-L (ROU.) [17] which is based on textual overlap, and the F1 of the BERTScore (BER.) [38] using the recommended microsoft/deberta-xlarge-mnli model, which is based on token-level semantic similarity.
@@ 4.3 Feedback Generation
We now show that we can improve both the correctness and alignment of generated feedback using our framework. We primarily focus on using Llama 2 Chat, an open-source LLM with 7B parameters, where we compare several versions of the model: Zero-Shot, i.e., simply prompting the base LLM, SFT, i.e., fine-tuning the base LLM on the gold feedback set, DPO (Score), i.e., training the LLM with DPO only on the augmented dataset, and DPO (Score + Mismatch), i.e., training the LLM with DPO on the augmented dataset and mismatched feedback. We additionally compare with the gold, human-written feedback in the dataset, as well as feedback generated by GPT-4. We use the same prompt for all methods, where we instruct the model to generate short and helpful feedback and to follow a version of the evaluation rubric.
@@ 4.3.1 Quantitative Analysis
Table 2 shows the average rubric labels and scores assigned by GPT-4 on all feedback in the test set, as well as the ROUGE-L and BERTScore values for reference. We see that DPO (Score + Mismatch) significantly improves the feedback scores compared to baselines (a 45% increase compared to Zero-Shot and SFT), showing that our data augmentation and training setup is highly effective at improving the quality of feedback generation with Llama 2. We additionally observe that including the mismatched feedback messages substantially increases the correctness of generated feedback, confirming their effectiveness as hard negative examples. Surprisingly, SFT does not outperform Zero-Shot on score, which shows that the standard fine-tuning setup is not effective for feedback generation. We can also see that ROUGE-L and BERTScore are unreliable estimates of feedback quality since they are highest on SFT, primarily because it copies the style of the gold feedback the closest.
@@
4.3.2 Qualitative Analysis
We also performed qualitative studies to compare the outputs of the different methods and find cases where they succeed or fail; the main findings are: 1) Zero-Shot produces feedback with the right style but struggles to follow instructions, particularly by not identifying the error or revealing the correct answer, and is prone to hallucinations and numerical errors. 2) SFT produces feedback that is generally short and blunt, and usually attempts to identify the error although is often incorrect. 3) DPO (Score) produces feedback that attempts to identify the error, adding details and questions to provide implicit suggestions and increase positivity. It also produces incorrect outputs although less so than SFT. 4) DPO (Score + Mismatch) is more accurate than DPO (Score) in identifying the error. 5) GPT-4 produces feedback with smooth and coherent language but tends to avoid mistakes by not clearly pointing out the error. When it does detail the error, it is often incorrect, although less often than the other methods. 6) For all methods, most failures occur because the LLM cannot correctly identify the error in the incorrect answer. This failure happens more often when multiple reasoning steps are required to arrive at the student’s answer. This observation suggests that although state-of-the-art LLMs claim to excel at mathematical reasoning, they cannot conceive how students may make errors [11, 18, 20]. One solution could be prompting the LLM to evaluate relevant calculations via code execution before writing the feedback [5, 31].
To provide more specific examples of several of our findings, we show a representative question and incorrect answer along with corresponding feedback generated by each method in Table 3. In this example, we see that the human-written feedback directly points out the error in the incorrect answer that the student only considered the denominators, and also suggests simplifying the fraction on the right. GPT-4 prefers to be vague to avoid mistakes by pointing out that the ratio should be simplified and gives a suggestion based on this, but does not explicitly identify why the student’s answer was wrong. Zero-Shot does not point out the student’s error and directly answers the question, even though it has been instructed to not do so; it then hallucinates an incorrect and irrelevant multiplication before finishing with a positive message. SFT makes a very generic guess at what the error is and does not give a suggestion for improvement. DPO (Score) guesses a more specific but incorrect error, gives a somewhat misleading suggestion, and uses questions to encourage the student. Finally, DPO (Score + Mismatch) correctly identifies the error and gives a helpful suggestion.
@@ 4.3.2 Qualitative Analysis
Human Not quite, the denominator has doubled but the fractions don’t have that relationship. Try simplifying 6/8 and then compare it to 1/4
GPT-4 Good effort, but remember to simplify the ratios first. Erica’s share is not twice as much, but a fraction of Rebecca’s. Try to compare 1/4 to 3/4 again. Keep going! ✓
Zero-shot Great effort! You are on the right track. Since the ratio is 1/4: 3/4, Erica has 1/4, which is a third of 3/4. To find the number of marbles Erica has, you can multiply 3/4 by 3, which gives you 9. So, Erica has 9 marbles. Keep practicing, and you’ll get it right next time!
SFT I think you have simplified the second fraction incorrectly.
DPO (Score) I think you have simplified 6/8 to 1/2. If we simplify 6/8 to 3/4, how many marbles does Erica have? How many does Rebecca have? What is the difference between them?
DPO (Score + Mismatch) I think you have found 8/4 = 2. But we need to simplify 6/8 to get the same units as 1/4. Then we can see how many more or less Erica has.
@@ 4.4 Feedback Evaluation
COR. 0.77 0.80 0.82 0.81
REV. 0.91 0.94 0.97 0.95
SUG. 0.73 0.76 0.73 0.73
DIA. 0.68 0.56 0.85 0.68
POS. 0.72 0.78 0.28 0.41
Avg. 0.76 0.77 0.73 0.71
Since we use GPT-4 to quantitatively evaluate feedback, we need to verify that GPT-4 can indeed label feedback accurately using our rubric. To do so, we randomly sample 80 feedback messages from the augmented test set and manually evaluate them on the rubric. We also recruited a graduate student with extensive teaching experience to evaluate these feedback messages for an additional set of human annotations. To reduce bias, we do not show human annotators GPT-4’s annotations or tell them whether a feedback message is human-written or LLM-generated. We compute agreement between human annotators using the Cohen’s kappa statistic, resulting in 0.53, 0.46, 0.35, 0.53 and 0.29 on the COR., REV., SUG., DIA. and POS. labels, respectively. This moderate agreement shows that while humans generally agree, they occasionally make errors due to the difficulty of the task, and have lower agreement on more subjective labels such as POS.
We measure the accuracy (Acc.), precision (Prec.), recall (Rec.), and F1 of GPT-4’s annotations with respect to human annotations on this set, and report the results averaged across both human annotators in Table 4. We observe that GPT-4 generally agrees with human annotations, with an average accuracy of 76% across labels. We also compute the Pearson correlation coefficient of the final rubric scores between GPT-4 and human annotations, resulting in 0.56 on average, indicating moderate overall correlation. In contrast, ROUGE-L and BERTScore both have average correlations of 0.40 with human-annotated rubric scores. Not only are these correlations smaller, but they are biased upward since the human-written feedback messages, which generally have high rubric scores, automatically get ROUGE-L and BERTScore values of 1.
@@
COR. (FP) Question: a jacket is reduced in a sale by $20\%$. a customer has a loyalty card for a further $10\%$ off the price. what is the total percentage reduction in the price of the jacket?
@@
COR. (FN) Question: i am thinking of a number. i divide it by 0.5 and my answer is 10. what number am i thinking of?
@@
DIA. (FP) Question: which fraction is equivalent $\frac{3}{7}$